# UniMoT: Unified Molecule-Text Language Model with Discrete Token Representation

### v1

@@ Abstract
The remarkable success of Large Language Models (LLMs) across diverse tasks has driven the research community to extend their capabilities to molecular applications. However, most molecular LLMs employ adapter-based architectures that do not treat molecule and text modalities equally and lack a supervision signal for the molecule modality. To address these issues, we introduce UniMoT, a Unified Molecule-Text LLM adopting a tokenizer-based architecture that expands the vocabulary of LLM with molecule tokens. Specifically, we introduce a Vector Quantization-driven tokenizer that incorporates a Q-Former to bridge the modality gap between molecule and text. This tokenizer transforms molecules into sequences of molecule tokens with causal dependency, encapsulating high-level molecular and textual information. Equipped with this tokenizer, UniMoT can unify molecule and text modalities under a shared token representation and an autoregressive training paradigm, enabling it to interpret molecules as a foreign language and generate them as text. Following a four-stage training scheme, UniMoT emerges as a multi-modal generalist capable of performing both molecule-to-text and text-to-molecule tasks. Extensive experiments demonstrate that UniMoT achieves state-of-the-art performance across a wide range of molecule comprehension and generation tasks.
@@ 1 Introduction
The incredible capabilities of Large Language Models (LLMs) [5, 44] have led to their widespread use as versatile tools for completing diverse real-world tasks. This success has sparked interest in Multi-modal LLMs [59, 52], which aim to enhance LLMs by enabling them to process multi-modal inputs and outputs. Prior research efforts [26, 41, 12, 6, 33, 35, 25] have focused on adapting LLMs to molecular tasks, resulting in the development of molecular LLMs. These molecular LLMs can analyze molecule structures [35, 33, 6], address drug-related inquiries [26, 41], assist in synthesis and retrosynthesis planning [12], support drug design [12], and more.
@@
Prevalent molecular LLMs commonly employ adapter-based architectures, adopting either a linear projection [26, 41, 6] or a Q-Former [33, 25] as an adapter to translate molecule features into the semantic space of LLM, as illustrated in Figure 1(a) and Figure 1(b). Despite demonstrating initial capabilities in molecular comprehension and yielding promising results in molecule-to-text generation tasks, they still fall short in text-to-molecule generation tasks. The critical issue within these methods is their unequal treatment of molecules and text, resulting in a lack of supervision for the molecule modality. This limitation significantly constrains model capacity and effectiveness.
@@
Discretizing continuous molecule features into discrete molecule tokens offers a promising solution for conducting both molecule-to-text and text-to-molecule generation tasks. By treating tokens from different modalities equally, we can predict the next molecule or text token in an autoregressive manner. However, directly discretizing molecule features poses several challenges: (i) This approach results in long sequences, with lengths equivalent to the number of atoms in a batch. LLMs typically experience a quadratic increase in computational complexity with sequence length [46]. (ii) Molecule tokens derived from molecule features lack left-to-right causal dependency, which conflicts with the unidirectional attention mechanism in LLMs. (iii) Molecule features lack textual information, hindering effective molecule-text interactions and alignment.
@@
To this end, we present UniMoT, a Unified Molecule-Text LLM that adopts a tokenizer-based architecture, integrating molecule comprehension and generation, as depicted in Figure 1(c). A pivotal aspect of UniMoT’s architecture is the molecule tokenizer for transforming molecules into molecule tokens. We introduce a Vector Quantization-driven [45] tokenizer, which incorporates a Q-Former [23] to bridge the modality gap between molecules and text. Specifically, we incorporate causal masks for the queries, enabling the Q-Former to generate a causal sequence of queries compatible with the unidirectional attention in LLMs. The sequence of queries is subsequently quantized into a sequence of molecule tokens using a learnable codebook. The molecule tokens encapsulate high-level molecular and textual information, which are then aligned with the latent space of a generative model via an MLP adapter, enabling the generation of desired molecules.
@@
Pretrained LLMs can integrate the molecule tokenizer by treating molecule tokens as new words and constructing a molecule vocabulary through mapping the learned codebook. We adopt the unified discrete token representation for molecules and text, coupled with the unified next-token-prediction training paradigm of LLM. This unification of representation and training paradigm enhances LLMs’ ability to understand molecule-text interactions and alignment. UniMoT interprets molecules akin to understanding a foreign language, and generates them as if they were text. Following a four-stage training scheme, UniMoT serves as a multi-modal generalist capable of performing both molecule comprehension and generation tasks.
@@
  • We introduce a molecule tokenizer specifically designed for LLMs, enabling the tokenization of molecules into short sequences of molecule tokens with causal dependency. These tokens encapsulate high-level molecular and textual information and can be decoded into desired molecules during inference.
  • We present UniMoT, a unified molecule-text LLM that adopts a tokenizer-based architecture instead of traditional adapter-based architectures. UniMoT unifies the modalities of molecule and text under a shared token representation and an autoregressive training paradigm.
  • UniMoT exhibits remarkable capabilities in multi-modal comprehension and generation. Extensive experiments demonstrate that UniMoT achieves state-of-the-art performance across a wide spectrum of molecule comprehension tasks and molecule generation tasks.
@@ 2 Related Works
The recent emergence of Vision Large Language Models (VLLMs) [24, 23, 28] has catalyzed advancements in Molecular LLMs, which encompass both single modality and multi-modality approaches. In the single modality domain, researchers are exploring diverse molecule representations, such as 1D sequences like SMILES strings [47, 8, 17], 2D molecule graphs [15, 56], 3D geometric conformations [56, 32], and textual information from the literature [43, 2, 21]. In the multiple modalities domain, various innovative approaches are being employed. MolT5 [11], a T5-based [38] model, is designed for SMILES-to-text and text-to-SMILES translations. Other works, such as MoMu [39], MoleculeSTM [31], MolFM [34], and GIT-Mol [29], leverage cross-modal contrastive learning to align the representation spaces of molecules and text. Additionally, some studies use multi-modal learning architectures to develop molecular LLMs, which often adopt adapter-based architectures. For instance, InstructMol [6], GraphGPT [41], and DrugChat [26] employ a simple projection layer to map molecule features to LLM’s input space. MolCA [33] and 3D-MoLM [25] utilize a Q-Former [23] to bridge the modality gap between molecules and text. However, these methods do not treat molecule and text modalities equally and lack a supervision signal for the molecule modality, limiting model capacity and effectiveness.
@@
Vector Quantization (VQ) [13] is a widely used technique in generative models. VQ-VAE [45] converts an image into a set of discrete codes within a learnable discrete latent space by learning to reconstruct the original image. VQ-GAN [57] enhances the generation quality by leveraging adversarial and perceptual objectives. In the context of molecules, VQ has been effectively applied to quantize molecule features. For example, DGAE [4] introduces a VQ model specifically for molecules, where molecules are encoded into discrete latent codes. Mole-BERT [54] uses VQ to rethink the pre-training of GNNs for molecular tasks. IMoLD [60] proposes using VQ to enhance invariant molecule representations, and VQSynergy [51] demonstrates the use of VQ for drug discovery.
@@ 3 Method
Our objective is to leverage the reasoning and generation capabilities of LLMs to enhance the comprehension and generation of molecule and text data. To achieve this, we focus on representing these modalities uniformly within the token representation, utilizing the next-token-prediction training paradigm of LLMs. As illustrated in Figure 2, we introduce a molecule tokenizer (Section 3.1) designed to transform molecules into molecule tokens by learning to reconstruct the input molecule. The molecule sequence can then be concatenated with the text sequence to form a multi-modal sequence, which is subsequently fed into an LLM for autoregressive pretraining (Section 3.2), as illustrated in Figure 3. The LLM vocabulary is expanded with molecule tokens mapped from the learned codebook. We introduce a four-stage training scheme for UniMoT (Section 3.3) comprising Causal Q-Former pretraining, molecule tokenizer pretraining, unified molecule-text pretraining, and task-specific instruction tuning. UniMoT is capable of performing both molecular comprehension and generation tasks following the training scheme.
@@ 3.1 Molecule Tokenizer for LLMs
We represent the structural information of a molecule as a graph, denoted by $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of atoms and $|\mathcal{V}|=N$ is the number of atoms. The task of the molecule encoder is to extract molecule features that are context-aware and encompass diverse local neighborhood structural information. By employing a molecule encoder, we obtain molecule features $\mathbf{X}\in\mathbb{R}^{N\times F}$, where each atom feature contains context-aware structural information.
@@
We employ a Q-Former model introduced by BLIP-2 [23] to generate queries $\mathbf{Z}=\{\boldsymbol{z}_{i}\}_{i=1}^{M}\in\mathbb{R}^{M\times d}$ containing high-level molecular and textual information, where $M$ represents the number of queries and $d$ denotes the dimension of queries. Specifically, we incorporate causal masks into the queries, ensuring that they only interact with preceding queries. This ensures the sequence of queries maintains a causal dependency, aligning with the requirements of LLMs operating on text sequence. Details regarding the Causal Q-Former can be found in Appendix A.
@@
$s_{i}=\operatorname{argmin}_{j\in\{1,\cdots,K\}}\left\|\boldsymbol{z}_{i}-\boldsymbol{c}_{j}\right\|_{2},\quad\text{for}\quad i=1,2,\cdots,M.$ (1)
Intuitively, the query $\boldsymbol{z}_{i}$ is quantized to the closest neighbor $\boldsymbol{c}_{s_{i}}$ in the codebook. As the vector quantization process is non-differentiable, we adopt the straight-through estimator [3] to train the Causal Q-Former by copying the gradient from the molecule tokens to the queries, as shown in Figure 2. The resulting embeddings of molecule tokens, denoted as $\mathbf{C}=\{\boldsymbol{c}_{s_{i}}\}_{i=1}^{M}$, are subsequently utilized for reconstructing molecules.
@@
An MLP adapter $\psi$ needs to be trained to align the discrete latent space of molecule tokens with the continuous latent space of a molecular generative model for molecule reconstruction. This can be represented as $\mathbf{X}_{R}=\psi(\mathbf{C})$, where $\mathbf{X}_{R}$ denotes the embeddings for reconstruction. To achieve alignment, we minimize the Mean Squared Error (MSE) loss between $\mathbf{X}_{R}$ and the SMILES [50] embeddings $\mathbf{X}_{S}$ produced by the pretrained SMILES encoder. Subsequently, we can reconstruct the molecule from $\mathbf{X}_{R}$ using the pretrained SMILES decoder. The training loss of the tokenizer is expressed as follows:
$\mathcal{L}_{\text{Tokenizer}}=\left\|\mathbf{X}_{R}-\mathbf{X}_{S}\right\|_{2}^{2}+\frac{1}{M}\sum_{i=1}^{M}\left\|\operatorname{sg}\left[\boldsymbol{z}_{i}\right]-\boldsymbol{c}_{s_{i}}\right\|_{2}^{2}+\frac{\beta}{M}\sum_{i=1}^{M}\left\|\operatorname{sg}\left[\boldsymbol{c}_{s_{i}}\right]-\boldsymbol{z}_{i}\right\|_{2}^{2}.$ (2)
@@ 3.2 Unified Molecule-Text Language Model
(a) Molecule-to-Text Autoregression.
(b) Text-to-Molecule Autoregression.
@@
While structural information is embedded in molecule features and captured by the molecule tokens through the tokenizer, we also aim to incorporate sequential information of molecules for better comprehension. Therefore, we concatenate the molecule sequence $\{s_{i}\}_{i=1}^{M}$ with the SMILES [50] sequence and a prompt to form the multi-modal input sequence $\{u_{i}\}_{i=1}^{L}$, as illustrated in Figure 3(a). The text sequence of the corresponding molecule caption is used as the generation target.

### v2

# Unified Molecule-Text Language Model with Discrete Token Representation

@@ Abstract
The remarkable success of Large Language Models (LLMs) across diverse tasks has driven the research community to extend their capabilities to molecular applications. However, most molecular LLMs employ adapter-based architectures that fail to equally integrate molecule and text modalities and lack explicit supervision signals for the molecular modality. To address these issues, we introduce UniMoT, a Unified Molecule-Text LLM adopting a tokenizer-based architecture that expands the vocabulary of LLMs with molecule tokens. Specifically, we introduce a Vector Quantization-driven tokenizer that incorporates a Q-Former to bridge the modality gap between molecule and text. This tokenizer transforms molecular structures into sequences of tokens exhibiting causal dependency, thereby encapsulating both high-level molecular features and textual information. Equipped with this tokenizer, UniMoT unifies molecule and text modalities under a shared token representation and an autoregressive training paradigm. This enables the model to process molecular structures as a distinct linguistic system and generate them in textual form. Through a four-stage training scheme, UniMoT functions as a multi-modal generalist capable of performing both molecule-to-text and text-to-molecule tasks. Extensive experiments demonstrate that UniMoT achieves state-of-the-art performance across a wide range of molecule comprehension and generation tasks.
@@ 1 Introduction
The incredible capabilities of Large Language Models (LLMs) Brown et al. (2020); Touvron et al. (2023) have led to their widespread use as versatile tools for completing diverse real-world tasks. This success has sparked interest in Multi-modal LLMs Zhan et al. (2024), which aim to enhance LLMs by enabling them to process multi-modal inputs and outputs. In fields like molecular science, Multi-modal LLMs present new opportunities by seamlessly integrating molecular data with textual information, opening up fresh possibilities for more efficient and accurate research and development.Prior research efforts Liang et al. (2023); Fang et al. (2023); Cao et al. (2023); Liu et al. (2023b); Li et al. (2024) have focused on adapting LLMs to molecular tasks, resulting in the development of molecular LLMs. These molecular LLMs can analyze molecule structures Liu et al. (2023b); Cao et al. (2023), address drug-related inquiries Liang et al. (2023), assist in synthesis and retrosynthesis planning Fang et al. (2023), support drug design Fang et al. (2023), and more.
@@
Prevalent molecular LLMs often use adapter-based architectures, such as linear projection Liang et al. (2023); Cao et al. (2023) or Q-Former Liu et al. (2023b); Li et al. (2024), to map molecule features into the LLM’s semantic space (Figure 1(a), Figure 1(b)). While effective in molecular comprehension and molecule-to-text generation, these models struggle with text-to-molecule generation. This is due to the reliance on adapters that require LLMs to directly generate SMILES strings Weininger (1988), a text-based representation of molecular structures. These architectures depend on strong alignment between SMILES and text, but as shown in Figure 1(a) and Figure 1(b), molecule and text modalities are not treated equally, with insufficient supervision for the molecular side, making alignment difficult.
@@
Discretizing continuous molecule features into discrete molecule tokens offers a promising solution for conducting both molecule-to-text and text-to-molecule generation tasks. By treating tokens from different modalities equally, we can predict the next molecule or text token in an autoregressive manner. However, directly discretizing molecule features poses several challenges: (i) This approach results in long sequences, with lengths equivalent to the number of atoms in a batch; (ii) Molecule tokens derived from molecule features lack left-to-right causal dependency, which conflicts with the unidirectional attention mechanism in LLMs; (iii) Molecule features lack textual information, hindering effective molecule-text interactions and alignment.
@@
To this end, we present UniMoT, a Unified Molecule-Text LLM that adopts a tokenizer-based architecture, integrating molecule comprehension and generation, as depicted in Figure 1(c). A pivotal aspect of UniMoT’s architecture is the molecule tokenizer for transforming molecules into molecule tokens. We introduce a Vector Quantization-driven Van Den Oord et al. (2017) tokenizer, which incorporates a Q-Former Li et al. (2023) to bridge the modality gap between molecules and text. Specifically, we incorporate causal masks for the queries, enabling the Q-Former to generate a causal sequence of queries compatible with the unidirectional attention in LLMs. The sequence of queries is subsequently quantized into a sequence of molecule tokens using a learnable codebook. The molecule tokens encapsulate high-level molecular and textual information, which are then aligned with the latent space of a pretrained generative model via an MLP adapter.
@@
Pretrained LLMs can integrate the molecule tokenizer by treating molecule tokens as new words and constructing a molecule vocabulary through mapping the learned codebook. We adopt the unified discrete token representation for molecules and text, coupled with the unified next-token-prediction training paradigm of LLM. This unification of representation and training paradigm enables effective molecule-text interactions and alignment through molecule-to-text and text-to-molecule autoregressive pretraining. For molecule generation tasks, UniMoT generates molecule tokens in an autoregressive manner rather than producing SMILES strings, and these molecule tokens can then be decoded into molecules using the generative model.
@@
  • We introduce a molecule tokenizer specifically designed for LLMs, enabling the tokenization of molecules into short sequences of tokens with causal dependency. These tokens encapsulate high-level molecular and textual information and can be decoded into desired molecules during inference.
  • We present UniMoT, a unified molecule-text LLM that adopts a tokenizer-based architecture instead of traditional adapter-based architectures. UniMoT unifies the modalities of molecule and text under a shared token representation and an autoregressive training paradigm. Following a four-stage training scheme, UniMoT effectively achieves molecule-text alignment.
  • UniMoT exhibits remarkable capabilities in multi-modal comprehension and generation. Extensive experiments show that UniMoT achieves state-of-the-art performance across a wide range of comprehension and generation tasks, while also offering a new perspective on molecule generation.
@@ 2 Related Works
Multi-modal Large Language Models.
Multi-modal Large Language Models (LLMs): Current multi-modal LLMs are typically built on a pre-trained LLM backbone and can understand multiple modalities. LLaVA Liu et al. (2024a) connects the image encoder to the LLM using a simple linear projection, while BLIP-2 Li et al. (2023) extracts high-level features from images with CLIP Radford et al. (2021) and uses Q-Former to reduce image token counts. While these models excel at multi-modal comprehension, they often lack focus on multi-modal generation. To address this, recent work unifies multi-modal comprehension and generation, such as SEED-LLaMA Ge et al. (2023) and AnyGPT Zhan et al. (2024), which unify processing across different modalities. Inspired by these advances, we introduce a tokenizer-based architecture in the molecule-text domain, converting molecular features into tokens compatible with LLMs.
@@
The recent emergence of Vision Large Language Models (VLLMs) Li et al. (2023) has catalyzed advancements in molecular LLMs, which encompass both single modality and multi-modality approaches. In the single modality domain, researchers are exploring diverse molecule representations, such as 1D sequences like SMILES strings Irwin et al. (2022), 2D molecule graphs You et al. (2020), 3D geometric conformations You et al. (2020), and textual information from the literature Taylor et al. (2022). In the multiple modalities domain, various innovative approaches are being employed. MolT5 Edwards et al. (2022), a T5-based Raffel et al. (2020) model, is designed for SMILES-to-text and text-to-SMILES translations. Other works, such as MoMu Su et al. (2022), MoleculeSTM Liu et al. (2023a), and GIT-Mol Liu et al. (2024b), leverage cross-modal contrastive learning to align the representation spaces of molecules and text. Additionally, some studies Cao et al. (2023); Liang et al. (2023); Liu et al. (2023b); Li et al. (2024) use multi-modal learning architectures to develop molecular LLMs, which often adopt adapter-based architectures. However, these methods do not treat molecule and text modalities equally and lack a supervision signal for the molecule modality, limiting model capacity and effectiveness.
@@
Vector Quantization (VQ) Gray (1984) is a widely used technique in generative models. VQ-VAE Van Den Oord et al. (2017) converts an image into a set of discrete codes within a learnable discrete latent space by learning to reconstruct the original image. VQ-GAN Yu et al. (2021) enhances the generation quality by leveraging adversarial and perceptual objectives. In the context of molecules, VQ has been effectively applied to quantize molecule features. For example, DGAE Boget et al. (2023) introduces a VQ model specifically for molecules, where molecules are encoded into discrete latent codes. Mole-BERT Xia et al. (2022) uses VQ to rethink the pre-training of GNNs for molecular tasks. IMoLD Zhuang et al. (2024) proposes using VQ to enhance invariant molecule representations, and VQSynergy Wu et al. (2024) demonstrates the use of VQ for drug discovery.
@@ 3 Method
Our objective is to leverage the reasoning and generation capabilities of LLMs to enhance the comprehension and generation of molecule and text data. To achieve this, we focus on representing these modalities uniformly within the token representation, utilizing the next-token-prediction training paradigm of LLMs. As illustrated in Figure 2, we introduce a molecule tokenizer (Section 3.1) designed to transform molecules into molecule tokens by learning to reconstruct the input molecule. The molecule sequence can then be concatenated with the text sequence to form a multi-modal sequence, which is fed into an LLM for molecule-to-text and text-to-molecule autoregressive pretraining (Section 3.2), as illustrated in Figure 3. The LLM vocabulary is expanded with molecule tokens mapped from the learned codebook. We introduce a four-stage training scheme for UniMoT (Section 3.3) comprising Causal Q-Former pretraining, molecule tokenizer pretraining, unified molecule-text pretraining, and task-specific instruction tuning. UniMoT is capable of performing molecule comprehension and generation tasks following the training scheme.
(a) Molecule-to-Text Autoregression.
(b) Text-to-Molecule Autoregression.
@@ 3.1 Molecule Tokenizer for LLMs
We represent the structural information of a molecule as a graph, denoted by $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of atoms and $|\mathcal{V}|=N$ is the number of atoms. The task of the molecule encoder is to extract molecule features that are context-aware and encompass diverse local neighborhood structural information. By employing a molecule encoder, we obtain molecule features $\mathbf{X}\in\mathbb{R}^{N\times F}$, where $F$ denotes the dimensionality of the feature vector for each atom.
@@
We employ a Q-Former model introduced by BLIP-2 Li et al. (2023) to generate queries $\mathbf{Z}=\{\boldsymbol{z}_{i}\}_{i=1}^{M}\in\mathbb{R}^{M\times d}$ containing high-level molecular and textual information, where $M$ represents the number of queries and $d$ denotes the dimension of queries. The Q-Former operates as a query-based transformer that utilizes learnable queries $\{\boldsymbol{z}_{i}\}_{i=1}^{M}$ to interact with molecule features $\mathbf{X}$ extracted by the molecule encoder. Specifically, we incorporate causal masks into the queries, ensuring that they only interact with preceding queries. This ensures the sequence of queries maintains a causal dependency, aligning with the unidirectional requirements of LLMs operating on text sequence. Details regarding the Causal Q-Former can be found in Appendix A.
@@
$s_{i}=\operatorname{argmin}_{j\in\{1,\cdots,K\}}\left\|\boldsymbol{z}_{i}-\boldsymbol{c}_{j}\right\|_{2},\quad\text{for}\quad i=1,2,\cdots,M.$
Intuitively, the query $\boldsymbol{z}_{i}$ is quantized to the closest neighbor $\boldsymbol{c}_{s_{i}}$ in the codebook. As the vector quantization process is non-differentiable, we adopt the straight-through estimator Bengio et al. (2013) to train the Causal Q-Former by copying the gradient from the molecule tokens to the queries, as shown in Figure 2. The resulting embeddings of molecule tokens $\{s_{i}\}_{i=1}^{M}$, denoted as $\mathbf{C}=\{\boldsymbol{c}_{s_{i}}\}_{i=1}^{M}$, are subsequently utilized for reconstructing molecules.
@@
An MLP adapter $\psi$ needs to be trained to align the discrete latent space of molecule tokens with the continuous latent space of a molecular generative model for molecule reconstruction. This can be represented as $\mathbf{X}_{R}=\psi(\mathbf{C})$, where $\mathbf{X}_{R}$ denotes the embeddings for reconstruction. To achieve alignment, we minimize the Mean Squared Error (MSE) loss between $\mathbf{X}_{R}$ and the SMILES Weininger (1988) embeddings $\mathbf{X}_{S}$ produced by the pretrained SMILES encoder. Subsequently, we can reconstruct the molecule from $\mathbf{X}_{R}$ using the pretrained SMILES decoder. The training loss of the tokenizer is expressed as follows:
$\mathcal{L}_{\text{Tokenizer}}=\left\|\mathbf{X}_{R}-\mathbf{X}_{S}\right\|_{2}^{2}+\frac{1}{M}\sum_{i=1}^{M}\left\|\operatorname{sg}\left[\boldsymbol{z}_{i}\right]-\boldsymbol{c}_{s_{i}}\right\|_{2}^{2}\\ +\frac{\beta}{M}\sum_{i=1}^{M}\left\|\operatorname{sg}\left[\boldsymbol{c}_{s_{i}}\right]-\boldsymbol{z}_{i}\right\|_{2}^{2}.$ (2)
@@ 3.2 Unified Molecule-Text Language Model
[…]
KV-PLM 70.50 72.12 55.03 59.83 89.17 54.63 65.40 78.50
AttrMask 67.79 75.00 63.57 58.05 75.44 73.76 75.44 80.28
InfoGraph 64.84 76.24 62.68 59.15 76.51 72.97 70.20 77.64
MolCLR 67.79 75.55 64.58 58.66 84.22 72.76 75.88 71.14
GraphMVP 68.11 77.06 65.11 60.64 84.46 74.38 77.74 80.48
MoleculeSTM 69.98 76.91 65.05 60.96 92.53 73.40 76.93 80.77
InstructMol (Vicuna-7B) 70.00 74.67 64.29 57.80 91.48 74.62 68.90 82.30
UniMoT (Llama-2-7B) 71.37 76.43 65.78 59.79 92.89 75.97 78.49 83.69
Table 1: ROC-AUC (%) of molecular property prediction task (classification) on the MoleculeNet Wu et al. (2018) datasets. Bold indicates the best performance and underline indicates the second best performance.
@@
While structural information is embedded in molecule features and captured by the molecule tokens through the tokenizer, we also aim to incorporate sequential information of molecules for better comprehension. Therefore, we concatenate the molecule sequence $\{s_{i}\}_{i=1}^{M}$ with the SMILES Weininger (1988) sequence and a prompt to form the multi-modal input sequence $\{u_{i}\}_{i=1}^{L}$, as illustrated in Figure 3(a). The corresponding molecule caption is used as the generation target.